OpenStudy (anonymous):

How do you simplify (-1+sqrt3) over (2-sqrt 3)?

OpenStudy (earthcitizen):

divide by the conjugate of the denominator

OpenStudy (anonymous):

\[\frac{-1+\sqrt{3}}{2-\sqrt{3}}*\frac{2+\sqrt{3}}{2+\sqrt{3}}\]

OpenStudy (earthcitizen):

\[(2+\sqrt{3})\] top and bottom

OpenStudy (anonymous):

Okay I think I wrote it wrong, it's actually 2 times sqrt 3 on the bottom, not 2 - sqrt 3. Sorry!

OpenStudy (anonymous):

Thank you so much for the replies!

OpenStudy (anonymous):

\[\frac{-1+\sqrt{3}}{2\sqrt{3}}*\frac{\sqrt{3}}{\sqrt{3}}\]

OpenStudy (anonymous):

You want a number that'll rationalize the denominator, \[(\sqrt{3})^2=3\]